# Hyperbolic Fracton Model, Subsystem Symmetry and Holography III: Extension to Generic Tessellations

### v1

@@ VI Rindler Reconstruction
The reconstruction begins by assigning values to all spins on the boundary subregion $A$. For each operator $\mathcal{O}_{v}=1$ adjacent to the boundary, if q1q-1 of its spins are known then the value of the remaining bulk spin is uniquely determined. By applying this procedure recursively, we can reconstruct the bulk spins inward until all constraints are exhausted. The reconstruction process halts when all remaining vertex operators involve two or more undetermined spins, making further reconstruction impossible. The reconstructed bulk region is enclosed by the minimal wedge, which is illustrated in 9.
@@ VII Mutual information
Having established the Rindler reconstruction, we now turn to the relationship between boundary entanglement and bulk geometry. In AdS/CFT, this is captured by the Ryu-Takayanagi (RT) formula, which relates entanglement entropy between two complimentary boundary subregions to the area of minimal covering surfaces in the bulk [38, 41]. In this section, we demonstrate that the HFM on arbitrary $\{p,q\}$ tesselation provides an analogue of this principle, where the mutual information between boundary subregions is determined by the length of minimal wedge.
@@ VIII Black Hole Entropy
The quantities $N^{i}_{\beta}$ and $\sum N^{k}_{\alpha}$ can be expressed analytically using Eqs. (11,12)
@@ IX.2 Scaling growth of fractons
Next, let us examine the cases when $p$ is odd. For the triangular tessellations, the number of fractons remains constant after the initial layer:
@@
$N_{\text{frac}}(5,4,k)=\begin{cases}10\times 3^{(k-2)/2}\quad\text{for even $k$}\\ 2\times 3^{(k+1)/2}\quad\text{for odd $k$}\end{cases}$ (36)
@@ X Conclusion
In this work, we have developed a comprehensive generalization of the hyperbolic fracton model (HFM) beyond the $\{5,4\}$ tessellation, establishing a unifying framework to understand fracton physics, holographic correspondence, and subsystem symmetry in negatively curved spaces. By systematically extending the model to generic $\{p,q\}$ tessellations, we revealed that the essential features of fracton physics, such as immobile excitations, subextensive ground-state degeneracy, and subsystem symmetries, persist in a wide class of hyperbolic geometries. However, the interplay between local constraints and geometric curvature give rise to qualitatively new phenomena, including layer-by-layer recursive degeneracy growth governed by the inflation rule and a new fracton propagation behavior.
@@
These results collectively establish that holography analogies in fracton models are not accidental but stem from a deeper geometric feature of subsystem symmetries and the combinatorial structure of hyperbolic space. Our framework provides a controlled setting in which generalized symmetries, curvature, and entanglement have intertwined relations. It opens the door to future investigations of quantum error correction, emergent gauge theories, and higher-rank tensor formulations on hyperbolic lattices. In particular, exploring the dynamical and finite-temperature behavior of the hyperbolic fracton phase, as well as its generalization to quantum fracton order, may illuminate new routes toward understanding the microscopic origin of holographic gravity in strongly correlated quantum matter, and provide new routes toward quantum error correction.

### v2

@@ V.4 Subsystem Symmetries
In particular, for all {3,q3,q} tessellations, the subsystem symmetry is extensive; any ground state is connected to an extensive number of other ground states via closed loop moves as illustrated for the {3,63,6} lattice in Fig. 8. This will have important consequences for the Rindler reconstruction discussed in the next chapter.
@@ VI Rindler Reconstruction
The reconstruction begins by assigning values to all spins on the boundary subregion $A$. For each operator $\mathcal{O}_{v}=1$ adjacent to the boundary, if q1q-1 of its spins are known then the value of the remaining bulk spin is uniquely determined. By applying this procedure recursively, we can reconstruct the bulk spins inward until all constraints are exhausted. The reconstruction process halts when all remaining vertex operators involve two or more undetermined spins, making further reconstruction impossible. The reconstructed bulk region is enclosed by the minimal wedge, which is illustrated in 11.
Essentially, the model needs to have sub-extensive subsystem symmetry to make Rindler reconstruction work. That is, the symmetry acts on the boundary and also extends in to the bulk infinitely deep in the thermodynamical limit. So whatever action applied to the boundary must be extended to the bulk within the ground state sector. Therefore, for $p=3$, the local symmetry which allows for an extensive number of loop moves does not support the Rindler reconstruction. This reconstruction does, however, work generally for $p\geq 4$.
@@ VII Mutual information
Having established the Rindler reconstruction, we now turn to the relationship between boundary entanglement and bulk geometry. In AdS/CFT, this is captured by the Ryu-Takayanagi (RT) formula, which relates entanglement entropy between two complimentary boundary subregions to the area of minimal covering surfaces in the bulk [37, 40]. In this section, we demonstrate that the HFM on arbitrary $\{p,q\}$ tessellation provides an analogue of this principle, where the mutual information between boundary subregions is determined by the length of minimal wedge.
@@ VIII Black Hole Entropy
The quantities $N^{i}_{\beta}$ and $\sum N^{k}_{\alpha}$ can be expressed analytically using Eqs. (12,13)
@@
Finally, we note that in this section we considered the “black hole” as a toy-model analogue rather than a gravitational black hole in the sense of general relativity. Our construction follows the spirit of holographic tensor-network models defined on discretized hyperbolic lattices [23], where a black hole is represented by removing a finite region in the internal of the lattice and introducing additional degrees of freedom associated with the resulting boundary. Because the lattice geometry is a regular $\{p,q\}$ tessellation of the hyperbolic plane, the curvature remains uniform everywhere and no metric singularity is present. The effective “horizon” therefore arises simply from the cutoff of a finite region of the lattice. The hyperbolic tiling terminate at this boundary and do not probe the interior region, with dangling degrees of freedom at the boundary, reproducing the black hole “soft hair” degrees of freedom. The full geometric and dynamical properties of gravitational black holes are not reproduced.
@@ IX.2 Scaling growth of fractons
Next, let us examine the cases when $p$ is odd. For the triangular tessellations, including the {3,6} Euclidean tessellation, the number of fractons remains constant after the initial layer:
@@
As a consequence, the two Euclidean tessellations with (sub-)extensive ground-state degeneracy preserve a constant number of fractons as they are expelled outwards.
@@
$N_{\text{frac}}(5,4,k)=\begin{cases}5\quad\text{for $k=1$}\\ 10\times 3^{(k-2)/2}\quad\text{for even $k$}\\ 2\times 3^{(k+1)/2}\quad\text{for odd $k>1$}\end{cases}$ (38)
@@ X Conclusion
In this work, we have developed a comprehensive generalization of the hyperbolic fracton model (HFM) beyond the $\{5,4\}$ tessellation, establishing a unifying framework to understand fracton physics, holographic correspondence, and subsystem symmetry in negatively curved spaces. By systematically extending the model to generic $\{p,q\}$ tessellations, we revealed that the essential features of fracton physics, such as immobile excitations, extensive ground-state degeneracy, and subsystem symmetries, persist in a wide class of hyperbolic geometries. The ground state degeneracy is finite and small for $q=3$, extensive for $q>3$ with the exception of {4,4} which is subextensive. In particular the ground state of the Euclidean {6,3} tessellation is also extensively degenerate. This suggests that the properties of the HFM are not only dictated by the negative curvature of the hyperbolic space, but also by the local connectivity $q$. In addition, for $p>4$ and $q>3$, the local structure of the tessellation enforces the number of fractons to grow exponentially in distance as they are expelled outwards, and algebraically in system size. Fractons of the HFM model are thus inherently different from type-I and type-II fractons on flat cubic lattices.
@@
These results collectively establish that holography analogies in fracton models are not accidental but stem from a deeper geometric feature of subsystem symmetries and the combinatorial structure of hyperbolic space. Our framework provides a controlled setting in which generalized symmetries, curvature, and mutual information have intertwined relations. It opens the door to future investigations of quantum error correction, emergent gauge theories, and higher-rank tensor formulations on hyperbolic lattices. In particular, quantum dynamics may be added to our model in a way analogue to how the quantum X-cube model has been extended to $H^{2}\times S^{1}$ hyperbolic space with an extra spatial dimension of $S^{1}$ and proper boundary conditions. Exploring the dynamical and finite-temperature behavior of the hyperbolic fracton phase, as well as its generalization to quantum fracton order, may illuminate new routes toward understanding the microscopic origin of holographic gravity in strongly correlated quantum matter, and toward quantum error correction.
Finally, we comment on the relation between the hyperbolic lattice studied here and holographic constructions motivated by the AdS/CFT correspondence. The geometry considered in this work is the hyperbolic plane $H^{2}$, which may be viewed as a spatial time slice of AdS3. However, it is important to emphasize that a simple vertical stacking of $H^{2}$ layers would instead produce the geometry $H^{2}\times\mathbb{R}$, which is not equivalent to AdS3. Our present construction therefore captures only the spatial structure of a hyperbolic slice, similar to many tensor-network realizations of holography defined on hyperbolic lattices. Incorporating a consistent dynamical time direction compatible with AdS geometry remains a nontrivial problem and lies beyond the scope of this work, but represents an interesting direction for future studies of fracton models in curved spacetimes.